# Mining-Gym: A Configurable RL Benchmarking Environment for Truck Dispatch Scheduling

### v1

@@ Abstract
Mining process optimization, particularly truck dispatch scheduling, is a critical factor in enhancing the efficiency of open-pit mining operations. However, the dynamic and stochastic nature of mining environments—characterized by uncertainties such as equipment failures, truck maintenance, and variable haul cycle times—poses significant challenges for traditional optimization methods. While Reinforcement Learning (RL) has demonstrated promise in adaptive decision-making for mining logistics, its practical deployment requires rigorous evaluation in realistic and customizable simulation environments. The lack of such standardized environment benchmarks limits fair algorithm comparisons, reproducibility, and real-world applicability of RL-based approaches in open-pit mining settings. To address this challenge, we introduce Mining-Gym, a configurable, open-source benchmarking environment designed for training, testing, and comparing RL algorithms in mining process optimization. Built on Discrete Event Simulation (DES) and seamlessly integrated with the OpenAI Gym interface, Mining-Gym offers a structured testbed that enables the direct application of advanced RL algorithms from Stable Baselines. The framework models key mining-specific uncertainties, such as equipment failures, queue congestion, and stochasticity of mining processes, ensuring a realistic and adaptive learning environment. Additionally, a graphic user interface (GUI) for easy parameter selection for mine-site configuration, comprehensive data logging system, a built-in KPI dashboard and real-time representative visualization of mine-site enables in-depth performance analysis, facilitating standardized, reproducible evaluation across multiple RL strategies and baseline heuristics.
Index Terms:
@@
I Introduction
@@
Mining process optimization aims to enhance efficiency and productivity by improving resource allocation, equipment scheduling, and material handling. However, these operations are highly complex, influenced by dynamic factors such as equipment failures, fluctuating ore quality, and unpredictable environmental conditions. Traditional optimization methods, such as linear programming and heuristics, struggle to adapt in real time, leading to inefficiencies and increased costs.
@@
RL offers a promising dynamic approach, but its adoption in mining remains limited due to the lack of standardized simulation environments that accurately model the stochastic nature of mining operations, provide a configurable testbed for algorithm development and comparison, and enable reproducible research with standardized scenarios and metrics. Additionally existing simulators are often proprietary, overly simplistic, or tailored to specific algorithms, hindering fair comparisons and slowing the transition from academic research to industrial application.
To address these, we introduce Mining-Gym, a configurable benchmarking framework for truck dispatch scheduling optimization. Unlike prior tools, Mining-Gym offers a comprehensive framework covering the following :
@@
  • A high-fidelity discrete-event simulation (DES) capturing real-world complexities,
@@
  • Seamless integration with OpenAI Gym for RL algorithm compatibility,
@@
  • Extensive customization for diverse mining scenarios,
  • Comprehensive logging and visualization tools, and
  • Open-source availability for collaborative research.
By standardizing benchmarking, Mining-Gym enables fair algorithm evaluation, facilitates reproducible experiments, and bridges the gap between theoretical RL development and industrial application.
The rest of this paper is structured as follows: Section II reviews related work in mining simulation and RL and higlights the limitations of conventional simulators. Section III details Mining-Gym’s architecture, integrating DES with RL. Section IV covers discussions on implementation features, including the configuration system and visualizations. Section V outlines experimental setup and Section VI presents and discusses the results. Finally section VII presents the conclusion and future work.
II Background and Related Work
II-A Background
II-A1 Truck dispatch scheduling optimization
Managing dispatch scheduling, particularly truck dispatching, represents a pivotal challenge in mining process optimization. Truck dispatching is dedicated to transporting extracted supply materials—both in quantity and quality—from mining fronts, where shovels excavate, to destinations like Crushers, or dumping sites. These dispatching decisions significantly impact operational efficiency, being critically important as a large portion of mining costs are linked to truck-shovel activities Truck dispatching tasks often utilize mathematical programming to reduce equipment waiting times and optimize production [1]. Whereas heuristic methods simplify decision-making by using practical experience instead of exhaustive optimization. These strategies include assigning trucks to the nearest shovel, prioritizing by equipment capacity or material demand, and leveraging historical data patterns [2]. Truck dispatching tasks often utilize static optimization techniques to minimize equipment waiting times and optimize production [1]. However, such conventional methods typically require re-optimization when complex configurations change, such as during equipment breakdowns, and may not be robust enough to handle the stochastic nature of minesite processes.
@@ II-A1 Truck dispatch scheduling optimization
Reference Year Uncertainties Simulator arch. RL adaptability RT Visualization Customizable Platform used Open source
@@
TABLE I: A study of conventional simulators used in Surface Mining process optimization, especially for Truck Dispatching. Comparative analysis of mining simulators based on key features. The table evaluates simulators across several dimensions: Uncertainties, ranging from Extensive (comprehensive modeling of equipment failures, maintenance, variable haul times) to Fair (modeling of limited uncertainties); Simulator Architecture, comparing (DES) and Simpler Rule-based Logic; RL Adaptability, which indicates whether the simulator is Adaptable (native integration), Custom (requires adaptation), or N.A. (not designed for RL); Real-time Visualization, measuring the extent of Extensive (Elaborate animated visuals with KPI dashboards), Fair (basic visual representation, or None.
However, due to the limitations of static optimization methods in dynamic environments, research is increasingly focusing on alternative, adaptive approaches.
II-A2 DES and integration with RL in Mining Operations
Recent research has explored combining RL with DES to simulate complex industrial environments. While standard RL environments like OpenAI Gym provide useful benchmarks, they often lack the realism required for industrial applications [17]. Industrial settings demand detailed, stochastic modeling, similar to DES, to account for dynamic conditions and resource constraints. Integrating RL’s trial-and-error learning with DES has gained traction for modeling real-world stochastic systems. For example, [18] transformed DES-based SCT controllers into RL environments, enhancing decision-making in automotive plant control.
II-A3 RL and the importance of simulator
RL excels in sequential decision-making, delivering state-of-the-art performance across various domains, including robotics, locomotion control, autonomous driving, and multi-agent systems. [21] The mining truck dispatching problem can be framed as a sequential decision-making task, where truck assignments must adapt to evolving conditions. RL provides a suitable framework for optimizing these dispatching strategies
II-A4 RL Applications in Truck Dispatch Optimization
II-A5 Limitations: Existing simulators
Refer to Table I for a comparison of conventional simulators based on crucial features. A notable absence of real-time visualization is observed in several studies, including [3, 5, 6, 7, 8, 9, 10, 13], which can hinder the ability to monitor and manage mining operations effectively. Customizability is another missing feature in [7], [9], and [10], limiting the flexibility of these simulators for adapting to different mining scenarios. Furthermore, the scarcity of open-source solutions, with only [12] providing such a framework, restricts broader adoption and collaborative improvement. A number of works, such as [7], [9], and [10], utilize rule-based architectures that may not be as adaptable or robust as more advanced simulation methods. Additionally, the proprietary nature of many conventional simulators limits research replication and comparison. Many also fail to adequately account for random events affecting key mining components like trucks, shovels, or crushers, as seen in [7]. Finally, many simulators are not designed for RL settings or are tailored to specific algorithms, limiting their versatility [8].
@@ II-A5 Limitations: Existing simulators
Figure 1: Overview of the Mining-Gym framework from different perspectives. (a) Event-driven or decision-point-based RL setting: The DES mining environment requests decisions from the RL policy only at specific decision points, such as when a truck requires a shovel assignment. (b) Comprehensive system architecture: Displays all key components of the Mining-Gym, including the graphical user interface (GUI), the generated configuration file used to initialize the environment, and the dashboard with real-time visualizations. (c) OpenAI Gym-compatible RL interface: Illustrates how the Mining-Gym integrates with OpenAI Gym by adapting simulation signals into the standard reset and step methods. This compatibility allows seamless integration with popular RL libraries, such as Stable-Baselines3 [25], enabling easy training and testing of RL models. Note: the reward function in (a) is for demonstration only.
III Mining-Gym: System Architecture and Design
The mining process especially the dynamics and uncertainty of the Load-Haul-Dump (crusher or dumping site)-Return-Query (LHDRQ) cycle is captured using a DES based simulation model. While the scheduling problem is modeled as a MDP in order to solve it using RL algorithms (see Fig.1). Following we discuss these two modeling processes.
III-A Mining-Gym overview
@@
The mining gym simulator and benchmarking tool is a sophisticated platform designed for RL applications in mining operations. It features a GUI-based interface where users input crucial parameter values that define the state of the mining site, equipment, and other relevant factors. These parameters are then translated into a human-readable and editable text file, allowing for easy adjustments directly within the simulation environment. Users can choose to either train a scheduler policy or run an existing one, which can be either classical or RL-based. The simulator includes an additional interface that supports the training and testing of modern RL algorithms, enabling advanced simulation capabilities. Furthermore, the tool provides a real-time dashboard displaying key performance indicators (KPIs) and a dynamic, visual representation of the mining site’s operations, offering users valuable insights into the system’s performance and the effects of various parameters on mining activities.
@@
We present a DES environment for mining processes, integrated with OpenAI Gym—rebranded as Gymnasium in 2023—for RL applications. Gymnasium is a widely used toolkit that standardizes interactions across diverse environments, from control tasks to robotics and video games. It enhances the original Gym with improved modularity, better support, and expanded features, enabling efficient testing, comparison, and debugging of RL algorithms. Our wrapper maps DES inputs/outputs to Gym’s reset and step methods, ensuring compatibility with RL frameworks like Stable Baselines. Stable Baselines provides well-tested implementations of RL algorithms such as PPO, A2C, and DQN, simplifying integration with Gym environments. The done flag signals episode termination, while info offers additional simulation insights. Together, Gymnasium and Stable Baselines create a powerful ecosystem for RL research, facilitating reproducible experimentation and benchmarking in complex domains like mining simulations.
III-B Mining Process Modeling: DES Overview
@@
For DES modeling we have used a comprehensive python package named Salabim [26], which offers process interaction methods, queue handling, resources, statistical sampling, and real-time 2D/3D animation capabilities.
@@
Figure 2: (A) Simplified mining system showing three key components: (1) Resource Handler managing resource availability and assignments, (2) Preemption Handler detecting breakdowns and managing repair processes (B) DES-RL interaction flow illustrating how the RL policy integrates with the DES. At decision points, the environment state is processed by the RL policy to determine resource assignments. Immediate or step rewards guide learning during simulation, while the episodic reward at shift (or episode) end updates the policy before environment reset. (C) Load-Haul-Dump-Return-Query (LHDRQ) cycle illustrating the truck’s journey through the mining process, which begins with querying the dispatcher for assignments, followed by loading material, hauling to the destination, dumping, and returning empty. Breakdown events, managed by the Preemption Handler, can interrupt operations at any stage.
In the SALABIM framework, the Truck component In the SALABIM framework, the Truck component follows a Load-Haul-Dump (crusher or dumping site)-Return-Query (LHDRQ) cycle. In the loading phase, a truck requests an available shovel as advised by the dispatcher and waits its turn. Once granted access, it undergoes loading, simulating material transfer time. The truck then enters the haul phase, traveling a predefined trajectory to the dump (crusher or dumping site), represented by holding for travel time. The dump choice is made at the loading site or shovel.
@@
At the dump, the truck enters the dump phase, requesting access to an available crusher or dumping site. After access, it performs dumping, mirroring the time needed to unload ore or waste. The return phase follows, where the truck travels back to the shovel, completing one haul cycle. The query phase then begins, where the truck seeks scheduling or resource allocation from the dispatcher. The Dispatcher agent is the core scheduling strategist, optimizing mining operations by balancing resource utilization and minimizing wait times. It employs basic strategies like fixed schedules or nearest-first, as well as advanced approaches such as RL-trained neural network policies.
A truck may request the following resource allocations: 1) Shovel, 2) Crusher, 3) Dumping site, and 4) Route (not currently considered). Mining-Gym’s dispatcher modules handle these requests (except routing) and supports baseline (e.g. Random) aswell as learned strategies (RL based). Only one dispatcher can be replaced with a learnable NN-based learnable policy at a time, with shovel allocation set as the default for all experiments.
Modeling breakdown events is crucial for simulating disruptions when trucks, shovels, or dumps become unavailable due to failures. These breakdowns impact loading, dumping, and transit, affecting efficiency and resource use. A preemption handler enhances realism by managing resource interruptions. For example, if a shovel breaks down mid-loading, the preemption handler pauses operations until repairs or replacement occur, accurately reflecting real-world maintenance disruptions.
Additional features : in the final Mining-Gym simulation build upon the basic workflow (Fig. 2).
@@
  • Dispatcher Strategies: Non-learnable, baseline dispatcher strategies (default random) for resources like Crushers and Dumping sites are included. These strategies allow for the evaluation of different dispatching approaches without relying on RL.
III-C Modeling the Dynamic Dispatching Problem as MDP
III-C1 Typical RL

### v2

# Mining-Gym: A Configurable RL Benchmarking Environment for Open-Pit Truck Dispatch Optimization

@@ Abstract
Optimizing the mining process—particularly truck dispatch scheduling—is a key driver of efficiency in open-pit operations. However, the dynamic and stochastic nature of these environments, marked by uncertainties such as equipment failures, truck maintenance, and variable haul cycle times, poses significant challenges for traditional optimization. While Reinforcement Learning (RL) shows strong potential for adaptive decision-making in mining logistics, practical deployment hinges on thorough evaluation in realistic, customizable simulation environments. Yet the absence of standardized benchmarking currently hampers fair algorithm comparison, reproducibility of results, and real-world applicability of RL-based solutions in open-pit mining. To bridge this gap, we present Mining-Gym—a configurable, open-source benchmarking environment for training, testing, and evaluating RL algorithms in mining process optimization. Mining-Gym is built on Salabim-based Discrete Event Simulation (DES) and integrates with Gymnasium (formerly OpenAI Gym), enabling seamless adoption of state-of-the-art RL algorithms from Stable Baselines3. The framework captures mining-specific uncertainties such as equipment failures, queue congestion, and stochastic behavior through an event-driven decision-point architecture. Mining-Gym provides a graphical user interface (GUI) for intuitive parameter configuration, extensive data logging, and real-time visualization—supporting thorough performance analysis and standardized, reproducible evaluation of diverse RL strategies and heuristic baselines. To validate benchmarking performance, we compare classical heuristics with RL-based scheduling across six scenarios ranging from normal operation to severe equipment failures. The results demonstrate that Mining-Gym is an effective, reproducible testbed for mining-logistics research—enabling fair evaluation of adaptive decision-making for real mines. The demonstrative experiments also show the strong performance potential of RL-trained schedulers, motivating further development in this direction.
keywords:
@@
[…]
[inst1]organization=School of Electrical Engineering and Robotics, Queensland University of Technology,addressline=2 George Street, city=Brisbane, postcode=4000, state=QLD, country=Australia
1 Introduction
@@
Mining process optimization aims to enhance efficiency and productivity by improving resource allocation, equipment scheduling, and material handling. However, these operations are inherently complex and uncertain, influenced by dynamic and stochastic factors such as equipment failures, fluctuating ore quality, and unpredictable environmental conditions. These characteristics introduce significant levels of imprecision and variability into decision-making, making it challenging for traditional optimization methods—such as linear programming and heuristics—to adapt effectively in real time. As a result, these conventional approaches often lead to inefficiencies and increased operational costs [1].
@@
Reinforcement Learning (RL), a branch of machine learning grounded in soft computing principles, offers a promising adaptive approach for handling such complexity. However, its adoption in mining remains limited, primarily due to the absence of standardized benchmarking environments that enable fair algorithmic comparison, reproducible experimentation, and systematic evaluation across diverse operational scenarios. Moreover, existing simulators are often proprietary [2], overly simplistic [3], or narrowly tailored to specific algorithms—factors that hinder fair comparisons and slow the transition from academic development to industrial application.
To bridge this gap, we present Mining-Gym, a configurable benchmarking framework for truck dispatch scheduling optimization, designed specifically to support standardized evaluation of RL and classical algorithms in mining contexts. Unlike prior tools, Mining-Gym provides a comprehensive and open-source platform that supports the development and evaluation of RL algorithms by incorporating:
  • 1.
@@
    A high-fidelity discrete-event simulation (DES) capturing real-world operational complexities and stochastic behaviors,
  • 2.
@@
    Seamless integration with OpenAI Gym to ensure compatibility with modern RL algorithms,
  • 3.
@@
    Extensive customization options to model diverse and realistic mining scenarios,
  • 4.
    A comprehensive system for logging, visualization, and performance tracking, and
  • 5.
    Open-source  [^1] Source code available at https://github.com/CBCodeHub/Mining-Gym availability to promote transparency, collaboration, and reproducibility in research.
By enabling standardized benchmarking and fostering robust comparisons across algorithms and heuristics, Mining-Gym supports the practical application of RL in mining logistics. It serves as a crucial bridge between theoretical research in soft computing and its deployment in real-world industrial systems. The framework’s primary contribution is providing researchers with a reliable testbed for rigorous algorithmic evaluation. As a proof-of-concept demonstration, we present comparative results between classical heuristics (random, fixed) and RL-based scheduling (PPO) across six operational scenarios, validating Mining-Gym’s effectiveness as a benchmarking platform.
The remainder of this paper is organized as follows: Section 2 reviews related work in mining simulation and reinforcement learning, emphasizing the limitations of current benchmarking approaches. Section 3 describes the architecture of Mining-Gym, which integrates discrete-event simulation with RL to facilitate standardized evaluation. Section 4 presents the implementation details, including the configuration system, visualization tools, and logging mechanisms that ensure reproducible experiments. Section 5 outlines the experimental setup and provides a demonstrative comparison between classical heuristics and RL-based scheduling to showcase the framework’s benchmarking capabilities. Section 6 discusses the results of these experiments. Finally, Section 7 concludes the paper and highlights directions for future work.
2 Background and Related Work
2.1 Background
2.1.1 Truck dispatch scheduling optimization
Open-pit mining extracts ore through a coordinated process: hydraulic shovels excavate ore and waste at the pit face, trucks transport material to processing facilities, crushers process the ore, and dumping sites stockpile waste. The dispatch problem focuses on optimizing truck-to-shovel assignments to minimize idle time while respecting equipment constraints. Key operational features in the mining setting are:
  • 1.
    Multi-destination routing: Ore is routed to crushers and waste to dumping sites; the dispatcher must balance this allocation.
  • 2.
    Equipment capacity constraints: Shovels and crushers have finite capacities, leading to potential queues and bottlenecks.
  • 3.
    Stochastic disruptions: Equipment failures and variable cycle times require adaptive, real-time decision-making.
  • 4.
    Cost drivers: Truck idle time dominates operational costs; minimizing queue congestion is therefore economically critical.
Dispatching decisions significantly influence operational efficiency, as a large portion of mining costs are linked to truck–shovel activities. Truck dispatching tasks often employ mathematical programming approaches to reduce equipment waiting times and optimize production [4].
Heuristic methods are widely used for their simplicity and flexibility, especially in dynamic mining environments where traditional optimization struggles with real-time disruptions like equipment failures or changing conditions [5]. Common strategies include assigning trucks to the nearest shovel, prioritizing by equipment capacity or material demand, and using historical data trends. However, they lack guarantees of global optimality and may underperform in complex or unfamiliar scenarios. In contrast, static optimization approaches are often applied to truck dispatching tasks to minimize equipment idle time and enhance production efficiency [4]. However, these traditional methods usually require full re-optimization when faced with complex changes—like equipment breakdowns—and are often not robust enough to manage the inherent uncertainty of mining operations.
Given the limitations of static optimization methods in handling dynamic and uncertain mining environments, there is a growing research emphasis on developing adaptive and flexible approaches that can respond effectively to real-time operational disruptions.
2.1.2 DES-RL integration in Mining Operations
Recent research has explored combining RL with DES to simulate complex industrial environments. While standard RL environments like OpenAI Gym provide useful benchmarks, they often lack the realism required for industrial applications [9]. Industrial settings demand detailed, stochastic modeling, similar to DES, to account for dynamic conditions and resource constraints. Integrating RL’s trial-and-error learning with DES has gained traction for modeling real-world stochastic systems. For example, [10] transformed DES-based SCT (supervisory control theory) controllers into RL environments, enhancing decision-making in automotive plant control.
2.1.3 RL and the importance of simulator
RL excels in sequential decision-making, delivering state-of-the-art performance across various domains, including robotics, locomotion control, autonomous driving, and multi-agent systems [16]. The mining truck dispatching problem can be framed as a sequential decision-making task, where truck assignments must adapt to evolving conditions. RL provides a suitable framework for optimizing these dispatching strategies
2.1.4 RL Applications in Truck Dispatch Optimization
@@
Ref. Year Uncertainties Simulator arch. RL adaptability RT Visualization Customizable Platform used OS*
@@
Table 1: Comparative analysis of mining simulators for truck dispatching based on key features. The table evaluates simulators across several dimensions: Uncertainties, ranging from Extensive (comprehensive modeling of equipment failures, maintenance, variable haul times) to Fair (modeling of limited uncertainties); Simulator Architecture, comparing (DES) and Simpler Rule-based Logic; RL Adaptability, which indicates whether the simulator is Adaptable (native integration), Custom (requires adaptation), or N.A. (not designed for RL); Real-time Visualization, measuring the extent of Extensive (Elaborate animated visuals with KPI dashboards), Fair (basic visual representation, or None. *OS stands for Open Source/ public availability of code or simulator for use.
2.1.5 Limitations: Existing simulators
Refer to Table 1 for a comparison of conventional simulators based on crucial features. A notable absence of real-time visualization is observed in several studies, including [22, 24, 13, 20, 14, 3, 21, 15], which can hinder the ability to monitor and manage mining operations effectively. Customizability is another missing feature in [20], [3], and [21], limiting the flexibility of these simulators for adapting to different mining scenarios. Furthermore, the scarcity of open-source solutions, with only [26] providing such a framework, restricts broader adoption and collaborative improvement. A number of works, such as [20], [3], and [21], utilize rule-based architectures that may not be as adaptable or robust as more advanced simulation methods. Additionally, the proprietary nature of many conventional simulators limits research replication and comparison. Many also fail to adequately account for random events affecting key mining components like trucks, shovels, or crushers, as seen in [20]. Finally, many simulators are not designed for RL settings or are tailored to specific algorithms, limiting their versatility [14].
@@ 2.1.5 Limitations: Existing simulators
Figure 1: (a) Comprehensive system architecture: Displays all key components of the Mining-Gym, including the graphical user interface (GUI), the generated configuration file used to initialize the environment, and the dashboard with real-time visualizations. (b) OpenAI Gym-compatible RL interface: Illustrates how the Mining-Gym integrates with OpenAI Gym by adapting simulation signals into the standard reset and step methods. This compatibility allows seamless integration with popular RL libraries, such as Stable-Baselines3 [27], enabling easy training and testing of RL models.
3 Mining-Gym: System Architecture and Design
Mining-Gym’s architecture integrates two complementary modeling approaches to tackle the truck dispatch scheduling challenge. It captures the dynamic, stochastic Load-Haul-Dump (crusher or dumping site)-Return-Queue (LHDRQ) cycle via a high-fidelity Discrete Event Simulation (DES) model, while the dispatch scheduling is formulated as a Markov Decision Process (MDP), enabling Reinforcement Learning (RL) for adaptive optimization. This dual-modeling framework represents both the operational complexity of mining and the sequential decision-making essential for intelligent dispatching. The following sections detail Mining-Gym’s architecture, covering the system overview, DES implementation, and MDP formulation for RL-based optimization.
3.1 Mining-Gym overview
@@
Mining-Gym is a mining gym simulator and benchmarking tool designed for RL applications in mining operations. It features a GUI-based interface where users input crucial parameter values that define the state of the mining site, equipment, and other relevant factors. These parameters are translated into a human-readable text file, where they are divided into fixed parameters (sampled once for the entire run, defining the infrastructure) and operational parameters (stochastically sampled per shift/episode, modeling daily variability). This structure allows for easy adjustments directly within the simulation environment, while separate scenario files are now used to manage failure conditions for clearer experimentation. Users can choose to either train a scheduler policy or run an existing one, which can be either classical or RL-based. The simulator includes an additional interface that supports the training and testing of modern RL algorithms, enabling advanced simulation capabilities. Furthermore, the tool provides a real-time dashboard displaying key performance indicators (KPIs) and a dynamic, visual representation of the mining site’s operations, offering users valuable insights into the system’s performance and the effects of various parameters on mining activities (see Fig. 1).
@@
We present a DES environment for mining processes, integrated with OpenAI Gym, (currently Gymnasium) —for RL applications. Gymnasium is a widely used toolkit that standardizes interactions across diverse environments, from control tasks to robotics and video games. It enhances the original Gym with improved modularity, better support, and expanded features, enabling efficient testing, comparison, and debugging of RL algorithms. Our wrapper maps DES inputs/outputs to Gym’s reset and step methods, ensuring compatibility with RL frameworks like Stable Baselines. Stable Baselines provides well-tested implementations of RL algorithms such as PPO, A2C, and DQN, simplifying integration with Gym environments. The done flag signals episode termination, while info offers additional simulation insights. Together, Gymnasium and Stable Baselines create a powerful ecosystem for RL research, facilitating reproducible experimentation and benchmarking in complex domains like mining simulations.
3.2 Mining Process Modeling: DES Overview
@@
For DES modeling, we have used a comprehensive python package named Salabim  [^2] Salabim is an open-source DES and animation library in Python. Available at https://www.salabim.org/ which offers process interaction methods, queue handling, resources, statistical sampling, and real-time 2D/3D animation capabilities.
@@
In the Salabim framework, the Truck component follows a Load–Haul– Dump–Return–Query (LHDRQ) cycle with stochastic parameters representing operational variability.
In the Loading phase, a truck requests an available shovel as advised by the Dispatcher (classical or RL based) and waits for its turn. Once granted access, it undergoes loading, with material transfer time sampled from a configurable probability distribution (see Table 8 for example parameterizations). The truck then enters the Haul phase, traveling a predefined trajectory to the dump (crusher or dumping site), represented by holding for travel time. Travel time depends on load state: loaded trucks travel at reduced speed due to fuel penalties, while empty returns occur at higher speed. Both speeds are sampled from configurable distributions to reflect operational variability. At the dump, the truck enters the Dump phase, requesting access to an available crusher or dumping site. After access, it performs dumping, with unloading time sampled from distributions that differ between crushers and dumping sites to reflect their operational characteristics. The Return phase follows, where the truck travels back to the shovel, completing one haul cycle. Stochastic disruptions can interrupt any phase: equipment failures occur according to Mean Time Between Failure (MTBF) distributions (typically modeled as Poisson processes), triggering the preemption handler that suspends operations until repairs complete after a Mean Time To Repair (MTTR) sampled from appropriate distributions. The Query phase then begins, where the truck seeks scheduling from the dispatcher. This event-driven approach ensures dispatch decisions account for truck capacity constraints, variable cycle times, and disruption-induced uncertainty. All distributional parameters are fully configurable through the Mining-Gym interface, allowing users to calibrate the simulation to specific mine characteristics or test algorithmic robustness across diverse operational settings.
@@
At the dump, the truck enters the Dump phase, requesting access to an available crusher or dumping site. After access, it performs dumping, mirroring the time needed to unload ore or waste. The return phase follows, where the truck travels back to the shovel, completing one haul cycle. The query phase then begins, where the truck seeks scheduling or resource allocation from the dispatcher. The Dispatcher agent is the core scheduling strategist, optimizing mining operations by balancing resource utilization and minimizing wait times. It employs basic strategies like fixed schedules or nearest-first, as well as advanced approaches such as RL-trained neural network policies. A truck may request the following resource allocations: 1) Shovel, 2) Crusher, 3) Dumping site, and 4) Route (not currently considered). Mining-Gym’s dispatcher modules handle these requests (except routing) and supports baseline (e.g. Random) aswell as learned strategies (RL based).
Figure 2: (A) Simplified mine-site simulation logic of Mining-Gym, showing three key components: (1) Resource Handler managing resource availability and assignments, (2) Preemption Handler detecting breakdowns and managing repair processes (B) Load-Haul-Dump-Return-Query (LHDRQ) cycle illustrating the truck’s journey through the mining process, which begins with querying the dispatcher for assignments, followed by loading material, hauling to the destination, dumping, and returning empty. Breakdown events, managed by the Preemption Handler, can interrupt operations at any stage. (C) DES-RL interaction flow illustrating how the RL policy integrates with the DES. At decision points, the environment state is processed by the RL policy to determine resource assignments. Immediate or step rewards guide learning during simulation, while the episodic reward at shift (or episode) end updates the policy before environment reset.
Algorithm 1 MineGym: Event-Driven DES-RL Synchronization
1:Simulation shift duration $T_{\text{shift}}$
2:Trajectories $\{(s_{t},a_{t},r_{t})\}_{t=0}^{T_{\text{shift}}}$
3:Initialize DES $\mathcal{S}_{\text{DES}}\leftarrow\texttt{Salabim.Environment()}$
4:Initialize RL Agent $\mathcal{A}_{\text{RL}}\leftarrow\texttt{Gym.Agent()}$
[…]
6:while $t_{\text{current}}<T_{\text{shift}}$ do
7:  // Decision Point Trigger (DES Pauses)
8:  $\texttt{event}\leftarrow\mathcal{S}_{\text{DES}}.\texttt{run\_until\_next\_decision\_event()}$
9:  $\tau\leftarrow\mathcal{S}_{\text{DES}}.\texttt{event\_time\_advance}$
10:  $d\leftarrow\texttt{event.\text{triggering\_entity}}$ (e.g., Truck requesting assignment)
11:  $\triangleright$ The DES execution is instantaneously paused at $t=\mathcal{S}_{\text{DES}}.\texttt{now()}$
12:  if $d.\texttt{is\_initial\_state}$ then
13:    $a_{t}\leftarrow\pi_{\text{classic}}(\text{shovel states})$ $\triangleright$ Use default policy for initialization
14:  else
15:    // RL Observation
16:    $s_{t}\leftarrow\mathcal{S}_{\text{DES}}.\texttt{get\_instantaneous\_state}(d)$ $\triangleright$ Snapshot of system state at exact moment of pause
17:    // Synchronous Communication (Gym Step)
18:    $\left(a_{t},r_{t}\right)\leftarrow\mathcal{A}_{\text{RL}}.\texttt{step\_and\_block}(s_{t},\tau)$ $\triangleright$ Agent computes action $a_{t}$, DES thread blocks until action received
19:    $\triangleright$ Reward $r_{t}$ is computed by the Agent, often using $\tau$
20:  end if
21:  // DES Resumes
22:  $d.\texttt{assign\_resource}(a_{t})$ $\triangleright$ Apply the chosen action to the truck entity
23:  $\mathcal{S}_{\text{DES}}.\texttt{resume\_and\_advance\_time}(\tau)$
24:  $t_{\text{current}}\leftarrow t_{\text{current}}+\tau$
25:end while
Additional features: The Mining-Gym workflow integrates DES modeling with RL-based decision-making through several key mechanisms illustrated in Fig. 2 and Algorithm 1.
Core System Architecture (Fig. 2(A)): The simulation operates through two critical handlers: (1) the Resource Handler manages resource availability, queuing, and assignments for shovels, crushers, and dumps; (2) the Preemption Handler detects equipment breakdowns and coordinates repair processes, ensuring realistic disruption modeling across all equipment types.
Truck Operation Cycle (Fig. 2(B)): Trucks follow the Load-Haul-Dump-Return- Query (LHDRQ) cycle, where each phase—from loading at shovels through hauling to dumps/crushers and returning—can be interrupted by breakdown events managed by the Preemption Handler. This event-driven architecture captures operational uncertainties inherent in mining logistics.
DES-RL Synchronization (Fig. 2(A) & Algorithm 1): At decision points (when trucks query for assignments), the DES pauses and transfers the current system state to the RL policy (Algorithm 1, lines 6-14). The policy computes resource assignments while the DES thread blocks (line 16), ensuring synchronous communication. Immediate rewards guide step-wise learning, while episodic rewards at shift end enable policy updates before environment reset (lines 17, 22). The complete workflow integrates several advanced features designed to enhance realism and flexibility, which are presented as follows.
  • 1.
@@
  • 4.
    Dispatcher Strategies: Non-learnable baseline dispatching strategies, including the default random policy, are incorporated. These baselines enable systematic evaluation of different dispatching approaches and provide a reference for comparing the performance of RL-based policies
For detailed modeling assumptions underlying our DES implementation, please refer to Appendix A.
3.3 Modeling the Dynamic Dispatching Problem as MDP
3.3.1 Typical RL